OpenStudy (anonymous):

Please someone help me with this....What is the solution of the system of equations? 3x+2y+z=7 5x+5y+4z=3 3x+2y+3z=1

OpenStudy (joannablackwelder):

Try subtracting the third equation from the first.

OpenStudy (anonymous):

I actually figured out already but thanks for trying to help :)

OpenStudy (joannablackwelder):

Great! I'm glad you didn't need me. :) Keep up the good work!

OpenStudy (jmark):

Given 3x+2y+z=7.............(1) 5x+5y+4z=3............(2) 3x+2y+3z=1 …..........(3) subtract equation (1) and (3) =>3x+2y+z=7 =>3x+2y+3z=1 ….-.....-....-.....- ..------------------ ...0+0-2z=6 =>z=-3 substitute z value in equation (1) and (2) =>3x+2y=10.....(1a) =>5x+5y=15.....(2a) divide equation (2a) by 5 =>x+y=3.......(2b)*2 =>2x+2y=6.....(2b) subtract (2b) and (1a) =>3x+2y=10 =>2x+2y=6 ….-....-......- .---------------- ...x=4 and 4+y=3 =>y=-1 hence x=4,y=-1 and z=-3 CBSE Previous Year Question Papers Class 10 http://goo.gl/HzqHts